A reversible heat pump operating between thermal reservoirs at 15°C and 10°C consumes 240 J of energy each second. How much heat energy is "pumped" into the house each second?
Answer #1

Since it's reversible, entropy increase of universe = 0. The "universe" here is the heat reservoirs at T1 and T2 and the working substance:

1st law: Q1 = Q2 + W
Heat pumped = Q2 at T2. T1 = room temp.

For zero entropy change, 2nd law gives Q2/T2 = Q1/T1 or Q2 = Q1T2/T1
So Q1 = Q1T2/T1 + W
Q1(1 - T2/T1) = W
Q1 = W/(1 - T2/T1) = 240/(1 - 283/288) = 13824J answer

(COP = Q1/W = 57.6)